(Corrected Trial Balance) The following trial balance of Blues Traveler Corporation does not balance.

Blues Traveler Corporation
Trial Balance
April 30, 2020

  Debit  Credit
Cash$  5,912
Accounts Receivable5,240
Supplies2,967
Equipment6,100
Accounts Payable
$  7,044
Common Stock
8,000
Retained Earnings
2,000
Service Revenue
5,200
Office Expense   4,320000000

$24,539$22,244

An examination of the ledger shows these errors.

1.    Cash received from a customer on the account was recorded (both debit and credit) as $1,380 instead of $1,830.

2.    The purchase on account of a computer costing $3,200 was recorded as a debit to Office Expense and a credit to Accounts Payable.

3.    Services were performed on account for a client, $2,250, for which Accounts Receivable was debited $2,250 and Service Revenue was credited $225.

4.    A payment of $95 for telephone charges was entered as a debit to Office Expense and a debit to Cash.

5.    The Service Revenue account was totaled at $5,200 instead of $5,280.


Instructions

From this information prepare a corrected trial balance, Please give an explanation
Answer #1

co Answor for corrected trial balance as of april 30. Evoy cash received from cestomer on account was, (1830-1380 = 450) . Thc. comportes debit value = $3.000 The inic CUC LS 001 office is reduced to 3200. The accounts pagoble is concelly eredited. iThe cash to be credited twice (from ErYONO => cash=636] cash - 6362-95-95 = 6362-190 - 6172 (cash reduced by por cash #6142 Ebebit credit 61:13 cash accounts recivable Supplies Equipment Accounts pagabbc common stock Retained Earning $4790 3.0967 $.4